# DeepWriter: A Fact-Grounded Multimodal Writing Assistant Based On Offline Knowledge Base

### v1

@@ Abstract
Large Language Models (LLMs) have demonstrated remarkable capabilities in various applications. However, their use as writing assistants in specialized domains like finance, medicine, and law is often hampered by a lack of deep domain-specific knowledge and a tendency to hallucinate. Existing solutions, such as Retrieval-Augmented Generation (RAG), can suffer from inconsistency across multiple retrieval steps, while online search-based methods often degrade quality due to unreliable web content. To address these challenges, we introduce DeepWriter, a customizable, multimodal, long-form writing assistant that operates on a curated, offline knowledge base. DeepWriter leverages a novel pipeline that involves task decomposition, outline generation, multimodal retrieval, and section-by-section composition with reflection. By deeply mining information from a structured corpus and incorporating both textual and visual elements, DeepWriter generates coherent, factually grounded, and professional-grade documents. We also propose a hierarchical knowledge representation to enhance retrieval efficiency and accuracy. Our experiments on financial report generation demonstrate that DeepWriter produces high-quality, verifiable articles that surpasses existing baselines in factual accuracy and generated content quality.
DeepWriter: A Fact-Grounded Multimodal Writing Assistant Based On Offline Knowledge Base
Song Mao1, Lejun Cheng1,2, Pinlong Cai1, Guohang Yan1, Ding Wang1, Botian Shi1, 1Shanghai AI Lab, 2Peking University Correspondence: caipinlong@pjlab.org.cn
1 Introduction
@@
The rapid advancement of Large Language Models (LLMs) and Multimodal Large Language Models (MLLMs) has unlocked a wide array of applications, from programming and education to complex task automation. Commercial models like Gemini DeepMind (2025), Claude 3.5 Anthropic (2024) and GPT-4o OpenAI (2025) have shown proficiency in leveraging web search and deep reasoning to assist users with real-world problems.
@@
In this paper, we investigate the application of LLMs as sophisticated writing assistants. While the context windows of modern LLMs are sufficient for processing entire novels Team et al. (2024); MiniMax et al. (2025); Yang et al. (2025), their application in professional writing often remains at a superficial level of data processing and summarization Wu et al. (2025). In specialized domains such as finance, medicine, and law, LLMs frequently fail to generate expert-level responses, and are prone to factual hallucinations. Two primary paradigms have emerged to mitigate these issues. The first, Retrieval-Augmented Generation (RAG) Lewis et al. (2020), connects LLMs to external knowledge bases. However, RAG can struggle with maintaining coherence and consistency across long-form documents, as multi-turn retrieval may introduce disjointed information Gu et al. (2025). The second paradigm involves using online search to gather relevant information Chen et al. (2024); Sun et al. (2025), but the variable quality of web content can compromise the overall quality and reliability of the generated document.
Figure 1: The overall framework of DeepWriter.
@@
This foundation allows DeepWriter to structurally organize the retrieved information for multimodal generation. Where standard RAG systems struggle to place visual elements effectively, we introduce a novel multimodal reranking system to intelligently determine the most appropriate location for images and charts within the text. Furthermore, to ensure the final output is fully verifiable and trustworthy, DeepWriter excels at generating and citing multimodal content with fine-grained granularity. Instead of citing an entire source file, which is a common practice that makes validation intractable for large documents, our system tracks and attributes information at the paragraph level. This approach guarantees that the generated document is not only coherent and professional but also factually grounded and reliable, directly addressing the core challenge of hallucination in LLMs.
We test DeepWriter on financial report generation task. Experiment results show that DeepWriter, with a simple yet effective framework, can achieve comparable performance with other leading open-source long-context models.
@@ 2 Related Work
Several projects have explored the integration of LLMs with public data sources for scientific question-answering and writing. Systems like OpenScholar Asai et al. (2024) and ScholarQA Singh et al. (2025) aim to assist researchers by surveying and synthesizing information. DeepResearcher trained a LLM-based deep research agents end-to-end through scaling reinforcement learning in real-word environments Zheng et al. (2025). DeepWriter extends this line of work by focusing on long-form, professional document generation from a controlled, offline corpus, with an emphasis on multimodal content.
@@
RAG is a powerful technique for grounding LLM outputs in external knowledge, addressing issues of outdated information and hallucination. Innovations like GraphRAG Edge et al. (2025), LightRAG Guo et al. (2024), and KAG Liang et al. (2024) have introduced more sophisticated retrieval mechanisms. However, generating long, coherent documents with RAG remains a challenge. DeepWriter’s structured pipeline and hierarchical knowledge representation are designed to improve coherence and factual consistency in long-form generation tasks.
@@
The concept of an AI writing assistant is popular, with a focus on improving human productivity by generating text in various styles. Frameworks like OmniThink have explored a "slow-thinking" approach to generate articles based on search engine results Xi et al. (2025). However, OmniThink did not address multimodal information or user customization. Other related systems like STORM Shao et al. (2024) and CO-STORM Jiang et al. (2024) focus on generating outlines and then "filling in the blanks" but rely on web searches, which can introduce noise. RAPID can generate comprehensive and knowledge-intensive articles based on RAG Gu et al. (2025). QRAFT is an LLM-based agentic framework that mimics the writing workflow of human fact-checkers Sahnan et al. (2025).  Xiong et al. (2025) proposed a general agent framework that achieves human like adaptive writing through recursive task decomposition and dynamic integration of three fundamental task types, i.e. retrieval, reasoning, and composition. DeepWriter differentiates itself by using a curated knowledge base and integrating multimodal content generation.
3 Methodology
3.1 Task Definition
Given a collected domain-specific corpus $\mathcal{K}=\{D_{1},\dots,D_{m}\}$ and a user-provided query or topic $Q$, our goal is to generate a long document $P$ with the following characteristics:
@@ 3.1 Task Definition
    Multimodal Content: The document should include relevant images, tables, and charts derived from the corpus.
@@
    Factual Grounding: All claims, figures, and critical viewpoints must be accompanied by citations pointing to their source in the corpus.
@@
3.2 System Architecture
The DeepWriter pipeline consists of an offline processing stage and an online generation stage, as illustrated in the Fig 1.
@@ 3.2 System Architecture
In the data processing stage, we aim to extract structured information from unstructured file formats such as PDF. We first use Fitz [^1] https://github.com/pymupdf/PyMuPDF or MinerU Wang et al. (2024) to extract text, tables, images from the original files. During the process, we keep all file and page metadata to avoid information loss when performing fact-checking. After extracting information from files, we choose appropriate chunk size to avoid chunks that are too small or too large. We also use advanced VLM such as Qwen2.5-VL Bai et al. (2025) to generate detailed image or table captions for multimodal retrieval.
We structure the extracted information into a hierarchical database. This involves a three-level representation (chunk-page-document) to facilitate efficient, multi-granularity retrieval. The top "document" level contains abstract concepts, such as year or domain information, the "page" level is an intermediate level, which trades offs the efficiency and granularity of retrieval, the final "chunk" level is composed of knowledge pieces, holding the raw source information. The data processing hierarchy is shown in Fig 2.
@@
2:  $Q^{\prime}\leftarrow\text{RewriteQuery}(Q)$
3:  $\mathcal{T}\leftarrow\text{DecomposeTask}(Q^{\prime})$
4:  $\mathcal{T}_{\text{multimodal}}\leftarrow\text{ExtractMultimodalTasks}(\mathcal{T})$
5:  Phase 2: Hierarchical Retrieval
6:  for each subtask $t_{i}\in\mathcal{T}$ do
7:     $\mathcal{R}_{i}\leftarrow\text{RetrieveFromKB}(t_{i},\mathcal{K})$
8:     $\mathcal{R}_{i}^{\text{text}}\leftarrow\text{FilterTextContent}(\mathcal{R}_{i})$
9:     $\mathcal{R}_{i}^{\text{visual}}\leftarrow\text{FilterVisualContent}(\mathcal{R}_{i})$
10:  end for
11:  Phase 3: Content Organization
12:  $\mathcal{S}\leftarrow\text{GenerateSectionTitles}(\mathcal{T})$
13:  $\mathcal{C}\leftarrow\text{ClusterBySections}(\mathcal{R}^{\text{text}},\mathcal{R}^{\text{visual}},\mathcal{S})$
14:  Phase 4: Section-by-Section Writing
15:  $P\leftarrow\emptyset$, $\mathcal{H}\leftarrow\emptyset$
16:  for each section $s_{j}\in\mathcal{S}$ do
17:     $\mathcal{C}_{j}\leftarrow\text{GetSectionContent}(\mathcal{C},s_{j})$
18:     $d_{j}\leftarrow\text{CreateDraft}(s_{j},\mathcal{C}_{j},\mathcal{H})$
19:     $d_{j}^{\prime}\leftarrow\text{RefineDraft}(d_{j},\mathcal{C}_{j})$
20:     $P\leftarrow P\cup\{d_{j}^{\prime}\}$
21:     $\mathcal{H}\leftarrow\mathcal{H}\cup\text{SummarizeSection}(d_{j}^{\prime})$
22:  end for
23:  Phase 5: Multimodal Integration
24:  $\mathcal{M}\leftarrow\text{CollectMultimodalContent}(\mathcal{R}^{\text{visual}})$
25:  for each visual element $m_{k}\in\mathcal{M}$ do
26:     $\text{score}_{k,\cdot}\leftarrow\text{ComputeRelevance}(m_{k},P)$
27:     $\text{pos}_{k}\leftarrow\text{OptimizePlacement}(\text{score}_{k,\cdot},P)$
28:     $P\leftarrow\text{InsertVisual}(P,m_{k},\text{pos}_{k})$
29:  end for
30:  Phase 6: Citation Generation
31:  for each claim $c_{l}$ in $P$ do
32:     $\text{source}_{l}\leftarrow\text{FindSource}(c_{l},\mathcal{R})$
33:     $\text{citation}_{l}\leftarrow\text{GenerateCitation}(\text{source}_{l})$
34:     $P\leftarrow\text{AddCitation}(P,c_{l},\text{citation}_{l})$
35:  end for
36:  Phase 7: Final Assembly
37:  $P\leftarrow\text{AddImagePaths}(P)$
38:  $P\leftarrow\text{ValidateCitations}(P)$
39:  $P\leftarrow\text{FormatDocument}(P)$
40:  return  $P$
@@
During the online stage, we aim to generate a long-form, multimodal, factual grounded article based on user query $Q$ and the knowledge base $\mathcal{B}$ generated in offline stage. The overall algorithm is introduced in Algorithm 1.
Firstly, the initial user query $Q$ is rewritten and expanded to better reflect the user’s intention, which mitigates ambiguity such as using abbreviations or general questions. After rewriting the user’s query, we decompose the task according to Fact, Data and Point. The fact contains new information or restricted information. For example, when talking about "What is the total trade volume of the world in 2024?", we need to specify the countries/regions that account for, which requires the agent to generate a subtask on which countries should be included. The data directly quantifies the performance, trend or severity, which connects the fact and provides a strong foundation for the point. The task decomposition module carefully asks questions to retrieve exact numerical values to improve reliability. Besides text data, there are data illustrated in tables, charts or figures. So, we also require the task decomposition module to design strategies for searching multimodal contents that are related to the topic. The Point are conclusion drawn by editors or authors, these conclusions are considered as a viewpoint as a reference.
Next, we perform retrieval according to decomposed sub-tasks or sub-queries based on the knowledge base. To enable unified multimodal retrieval, we utilize GME Zhang et al. (2025), a multimodal embedding model that supports images and texts as the inputs. We use cosine similarity as the metric to select top-$k$ relevant documents. The relevant documents are then clustered based on pre-generated section titles, which reduces information seeking cost when generating section contents.
Now, we perform the writing process for each section. We ask the agent to write section by section according to relevant documents, focusing on combining the facts, data and points and writing in a general-to-specific pattern, such that audiences can focus on different parts, achieving customization capability. To improve fluency, we require the agent to make a draft instead of writing directly. We also summarize the written part as a compressed context to prevent repeated generated content.
Finally, we put everything together. We integrate multimodal contents by compute the relevance among different modalities. Then, we add citations for each section, and these citations are fine-grained such that we can trace to their original paragraphs. We detail these two parts in the following two subsections.
3.3 Interleaved Image-Text Generation
@@
To address this challenge, we propose a two-stage approach: content relevance scoring and contextual placement optimization. First, we compute relevance scores between each multimodal element and every paragraph in the generated text using the same GME embedding model. This creates a relevance matrix that captures semantic similarity between visual and textual content.
@@ 3.4 Grounded Citation
4 Experiments
4.1 System Configuration
We use MinerU as PDF processing toolkit to process a bundle of PDFs. The preprocess data are stored in a Milvus database, besides the processed information, we also add the embedding generated via GME (gme-Qwen2-VL-2B-Instruct) Zhang et al. (2025) and a descriptive caption for images and tables via a advanced vision language model Qwen2.5-VL 7B Bai et al. (2025). We use Qwen2-7B Yang et al. (2024) as the LLM performing writing tasks.
4.2 Task & Benchmark
We evaluate the performance of DeepWriter on WTR dataset [^2] See https://www.wto.org/english/res_e/reser_e/wtr_e.htm for details. WTR (World Trade Report) contains the report produced by World Trade Organization, it covers the statistic details from 2001 to 2024. We download the 23 annual reports as the corpus $\mathcal{K}$ and ask DeepWriter to write a long article for a specific topic $Q$.
For evaluation, following Shao et al. (2024); Jiang et al. (2024); Xi et al. (2025), we adopt Prometheus2-7B Kim et al. (2024) as the judge model to evaluate the quality of the generate article from several dimensions including Interest Level; Coherence and Organization, Relevance and Focus; Coverage. The evaluation standard is given in Table 2.
4.3 Models
@@
    Long-Context MLLMs: We use state-of-the-art LLMs with long context windows to assess their zero-shot long-form writing capabilities. We use Qwen-Plus as the LLM to be compared in this paper, of which context window is 131,072 with maximum output of 8,192 tokens [^3] See the official website  https://www.alibabacloud.com/help/en/model-studio/what-is-qwen-llm for more details.. We also [TODO]
@@
5 Results
5.1 Performance on WTR dataset
@@
5.2 Case Study
@@
Success Case: Comprehensive Trade Analysis
We asked DeepWriter to generate a report on "Global Trade Patterns in Digital Services." The system successfully decomposed this complex topic into subtasks covering regional trade flows, sector-specific trends, and regulatory frameworks. The generated document included relevant charts from the WTR dataset showing trade volume trends and properly cited specific paragraphs from source documents.
The hierarchical knowledge representation enabled efficient retrieval of both high-level concepts and specific data points. The multimodal content placement algorithm correctly positioned trade flow diagrams near their corresponding textual descriptions, enhancing document readability. Fine-grained citations allowed readers to trace every claim back to specific source paragraphs.
Failure Case: Ambiguous Query Resolution
When given the query "Economic impact of recent policies," DeepWriter struggled due to the ambiguity of "recent policies" and the lack of temporal context in the corpus. The system generated a generic overview rather than focusing on specific policy impacts, resulting in lower relevance and coverage scores.
This case highlights the importance of query rewriting and the need for better temporal reasoning capabilities. The task decomposition module failed to identify the need for temporal constraints, leading to overly broad retrieval that diluted the document’s focus.
Analysis of Limitations
Our case studies reveal several areas for improvement. First, the system requires more sophisticated temporal reasoning to handle queries involving time-sensitive information. Second, the multimodal placement algorithm sometimes struggles with complex visual elements that span multiple topics. Finally, the citation system, while precise, can become verbose for documents with many small claims.
These insights inform our future work directions, particularly the development of more advanced query understanding and temporal reasoning capabilities.
6 Conclusion

### v2

@@ Abstract
Large Language Models (LLMs) have demonstrated remarkable capabilities in various applications. However, their use as writing assistants in specialized domains like finance, medicine, and law is often hampered by a lack of deep domain-specific knowledge and a tendency to hallucinate. Existing solutions, such as Retrieval-Augmented Generation (RAG), can suffer from inconsistency across multiple retrieval steps, while online search-based methods often degrade quality due to unreliable web content. To address these challenges, we introduce DeepWriter, a multimodal, long-form and fact-grounded writing assistant that operates on a curated, offline knowledge base. DeepWriter leverages a novel pipeline that involves task decomposition, outline generation, multimodal retrieval, and section-by-section composition with reflection. By deeply mining information from a structured corpus and incorporating both textual and visual elements, DeepWriter generates coherent, factually grounded, and professional-grade documents. To evaluate the performance of DeepWriter, we curate a benchmark containing five domains, experiment results on the curated benchmark demonstrate that DeepWriter produces high-quality, verifiable articles that surpasses existing baselines in factual accuracy and generated content quality.
Introduction
@@
The rapid advancement of Large Language Models (LLMs) and Multimodal Large Language Models (MLLMs) has unlocked a wide array of applications, from programming and education to complex task automation. Commercial models like Gemini (DeepMind 2025), Claude 3.5 (Anthropic 2024) and GPT-4o (OpenAI 2025) have shown proficiency in leveraging web search and deep reasoning to assist users with real-world problems.
@@
In this paper, we investigate the application of LLMs as sophisticated writing assistants. While the context windows of modern LLMs are sufficient for processing entire novels (Team et al. 2024; MiniMax et al. 2025; Yang et al. 2025), their application in professional writing often remains at a superficial level of data processing and summarization (Wu et al. 2025b). In specialized domains such as finance, medicine, and law, LLMs frequently fail to generate expert-level responses, and are prone to factual hallucinations. Two primary paradigms have emerged to mitigate these issues. The first, Retrieval-Augmented Generation (RAG) (Lewis et al. 2020), connects LLMs to external knowledge bases. However, RAG can struggle with maintaining coherence and consistency across long-form documents, as multi-turn retrieval may introduce disjointed information (Gu et al. 2025). The second paradigm involves using online search to gather relevant information (OpenAI 2025; Google Team 2025; xAI Team 2025; Perplexity Team 2024), but the variable quality of web content can compromise the overall quality and reliability of the generated document.
Figure 1: Illustration of DeepWriter. DeepWriter receives a query and a offline corpus to produce a long-form, multimodal and fact-grounded article.
@@
To integrate retrieved materials structurally, we propose a multi-stage pipeline, consisting of task decomposition, document clustering, section by section writing. Previous work have shown that such a multi-stage procedure can achieve superiror performance compared with single stage pipeline (Bai et al. 2025b; Wu et al. 2025a). By utilizing a multi-stage processing pipeline, we mitigate the overload of integrating external knowledge and internal knowledge.
Next, we believe that multimodal contents is more appealing to users compared to pure text reports, thus we propose a novel multimodal reranking system to intelligently determine the most appropriate location for images and charts within the text. This foundation allows DeepWriter to structurally organize the retrieved information for multimodal generation.
Finally, to ensure the final output is fully verifiable and trustworthy, DeepWriter excels at generating and citing multimodal content with fine-grained granularity. Instead of citing an entire source file, which is a common practice that makes validation intractable for large documents, our system tracks and attributes information at the paragraph level. This approach guarantees that the generated document is not only coherent and professional but also factually grounded and reliable, directly addressing the core challenge of hallucination in LLMs.
We test DeepWriter on report generation task corresponding to five domains. Experiment results show that DeepWriter, with a simple yet effective framework, can achieve comparable performance with other leading open-source long-context models.
@@
Several projects have explored the integration of LLMs with public data sources for scientific question-answering and writing. Systems like OpenScholar (Asai et al. 2024) and ScholarQA (Singh et al. 2025) aim to assist researchers by surveying and synthesizing information. DeepResearcher (Zheng et al. 2025) trained a LLM-based deep research agents end-to-end through scaling reinforcement learning in real-word environments. DeepWriter extends this line of work by focusing on long-form, professional document generation from a controlled, offline corpus, with an emphasis on multimodal content generation.
@@
RAG is a powerful technique for grounding LLM outputs in external knowledge, addressing issues of outdated information and hallucination. Innovations like GraphRAG (Edge et al. 2025), LightRAG (Guo et al. 2024), and KAG (Liang et al. 2024) have introduced more sophisticated retrieval mechanisms. However, generating long, coherent documents with RAG remains a challenge. DeepWriter’s structured pipeline and hierarchical knowledge representation are designed to improve coherence and factual consistency in long-form generation tasks.
@@
The concept of an AI writing assistant is popular, with a focus on improving human productivity by generating text in various styles. Frameworks like OmniThink (Xi et al. 2025) have explored a “slow-thinking” approach to generate articles based on search engine results. However, OmniThink did not address multimodal information and relies on web search capability. Other related systems like STORM (Shao et al. 2024) and CO-STORM (Jiang et al. 2024) focus on generating outlines and then “filling in the blanks” but rely on web searches, which can introduce noise, besides, the generated article are constrained in Wikipedia format. RAPID can generate comprehensive and knowledge-intensive articles based on RAG (Gu et al. 2025). QRAFT is an LLM-based agentic framework that mimics the writing workflow of human fact-checkers (Sahnan et al. 2025).  (Xiong et al. 2025) proposed a general agent framework that achieves human like adaptive writing through recursive task decomposition and dynamic integration of three fundamental task types, i.e. retrieval, reasoning, and composition. SuperWriter (Wu et al. 2025a) proposed a reflection-driven agent framework that decomposes long-form text generation into a planning, writing, and refining loop, enabling iterative self-improvement through hierarchical direct preference optimization. DeepWriter differentiates itself by using a curated knowledge base and integrating multimodal content generation.
Methodology
Task Definition
Our goal is to generate a long document $P$ given a collected domain-specific corpus $\mathcal{K}=\{D_{1},\dots,D_{m}\}$ and a user-provided query or topic $Q$, with the following characteristics:
@@
    Multimodal Content: The document should include relevant images, tables, and charts derived from the corpus to enrich the modality of the document.
@@
    Factual Grounding: All claims, figures, and critical viewpoints must be accompanied by citations pointing to their source in the corpus to minimizing the hallucination bought by the LLMs.
@@
System Architecture
The DeepWriter pipeline consists of two stages: an offline processing stage that is responsible for converting the unstructured PDFs into unified format for retrieval, and an online writing stage, responsible for generating the final output, the framework of DeepWriter is as illustrated in the Fig 2.
Figure 2: The overall framework of DeepWriter.
@@
In the data processing stage, we aim to extract structured information from unstructured file formats such as PDF. We first use Fitz [^1] https://github.com/pymupdf/PyMuPDF or MinerU (Wang et al. 2024) to extract text, tables, images from the original unstructured documents. During the process, we keep all file and page metadata to avoid information loss when performing fact-checking. After extracting information from files, we choose appropriate chunk size to avoid chunks that are too small or too large. We also use advanced VLM such as Qwen2.5-VL (Bai et al. 2025a) to generate detailed image or table captions for multimodal retrieval.
We structure the extracted information into a hierarchical database. This involves a three-level representation (chunk-page-document) to facilitate efficient, multi-granularity retrieval. The top “document” level contains abstract concepts, such as year or domain information, the “page” level is an intermediate level, which trades offs the efficiency and granularity of retrieval, the final “chunk” level is composed of knowledge pieces, holding the raw source information. The data processing hierarchy is shown in Fig 3.
@@
[…]
4:Phase 2: Hierarchical Retrieval
5:for each subtask $t_{i}\in\mathcal{T}$ do
6:  $\mathcal{R}_{i}\leftarrow\text{RetrieveFromKB}(t_{i},\mathcal{K})$
7:  $\mathcal{R}_{i}^{\text{text}}\leftarrow\text{FilterTextContent}(\mathcal{R}_{i})$
8:  $\mathcal{R}_{i}^{\text{visual}}\leftarrow\text{FilterVisualContent}(\mathcal{R}_{i})$
9:end for
10:Phase 3: Content Organization
[…]
13:Phase 4: Section-by-Section Writing
[…]
15:for each section $s_{j}\in\mathcal{S}$ do
16:  $\mathcal{C}_{j}\leftarrow\text{GetSectionContent}(\mathcal{C},s_{j})$
17:  $d_{j}\leftarrow\text{CreateDraft}(s_{j},\mathcal{C}_{j},\mathcal{H})$
18:  $d_{j}^{\prime}\leftarrow\text{RefineDraft}(d_{j},\mathcal{C}_{j})$
19:  $P\leftarrow P\cup\{d_{j}^{\prime}\}$
20:  $\mathcal{H}\leftarrow\mathcal{H}\cup\text{SummarizeSection}(d_{j}^{\prime})$
21:end for
22:Phase 5: Multimodal Integration
[…]
24:for each visual element $m_{k}\in\mathcal{M}$ do
25:  $\text{score}_{k,\cdot}\leftarrow\text{ComputeRelevance}(m_{k},P)$
26:  $\text{pos}_{k}\leftarrow\text{OptimizePlacement}(\text{score}_{k,\cdot},P)$
27:  $P\leftarrow\text{InsertVisual}(P,m_{k},\text{pos}_{k})$
28:end for
29:Phase 6: Citation Generation
30:for each claim $c_{l}$ in $P$ do
31:  $\text{source}_{l}\leftarrow\text{FindSource}(c_{l},\mathcal{R})$
32:  $\text{citation}_{l}\leftarrow\text{GenerateCitation}(\text{source}_{l})$
33:  $P\leftarrow\text{AddCitation}(P,c_{l},\text{citation}_{l})$
34:end for
35:Phase 7: Final Assembly
[…]
39:return  $P$
@@
During the online stage, we aim to generate a long-form, multimodal, factual grounded article based on user query $Q$ and the knowledge base $\mathcal{B}$ generated in offline stage. The overall algorithm is introduced in Algorithm 1. We detail our key design in the following subsections.
Multi-Stage generation pipeline
Firstly, the initial user query $Q$ is rewritten and expanded to better reflect the user’s intention, which mitigates ambiguity such as using abbreviations or general questions. After rewriting the user’s query, we decompose the task according to Fact, Data and Point. The fact contains new information or restricted information. For example, when talking about ”What is the total trade volume of the world in 2024?”, we need to specify the countries/regions that account for, which requires the agent to generate a sub-task on which countries should be included. The data directly quantifies the performance, trend or severity, which connects the fact and provides a strong foundation for the point. The task decomposition module carefully asks questions to retrieve exact numerical values to improve reliability. Besides text data, there are data illustrated in tables, charts or figures. So, we also require the task decomposition module to design strategies for searching multimodal contents that are related to the topic. The Point are conclusion drawn by editors or authors, these conclusions are considered as a viewpoint as a reference.
Next, we perform retrieval according to decomposed sub-tasks or sub-queries based on the knowledge base. To enable unified multimodal retrieval, we utilize a unified multimodal embedding model that supports images and texts as the inputs, this unified embedding model alleviates the burden of loading multiple model and improves the searching efficiency. We then use cosine similarity as the metric to select top-$k$ relevant documents for each decomposed query. The relevant documents are then clustered based on pre-generated section titles, which reduces information seeking cost when generating section contents.
Now, we perform the writing process for each section. We ask the agent to write section by section according to relevant documents, section title and content that is already written, focusing on combining the facts, data and points and writing in a general-to-specific pattern, such that audiences can focus on different parts, achieving customization capability. To improve fluency, we require the agent to make a draft instead of writing directly. After completing the wriitng process, we ask the agent to summarized the written content to compress the context and prevent from generating repeated content.
Finally, we put everything together, including contents of each section, the multimodel contents and corresponding references. We integrate multimodal contents by compute the relevance among different modalities. Then, we add citations for each section, and these citations are fine-grained such that we can trace to their original paragraphs.
Interleaved Image-Text Generation
@@
To address this challenge, we propose a two-stage approach: content relevance scoring and contextual placement optimization. First, we compute relevance scores between each multimodal element and every paragraph in the generated text using the same embedding model. This creates a relevance matrix that captures semantic similarity between visual and textual content.
@@
Experiments
System Configuration
We use MinerU (Wang et al. 2024) as PDF processing toolkit to process a bundle of PDFs. MinerU accepts multiple format files as input and output the OCR results in a markdown format. We store the extracted images and texts in a Milvus database, besides the processed information, we also add the embedding generated via GME (gme-Qwen2-VL-2B-Instruct [^2] See https://huggingface.co/Alibaba-NLP/gme-Qwen2-VL-2B-Instruct(Zhang et al. 2025) as the retrieval key. For image and text items, a descriptive caption is generated via an advanced vision language model Qwen2.5-VL 7B (Bai et al. 2025a). We use Qwen2-7B (Yang et al. 2024) as the LLM performing online generation tasks.
Evaluation Setting
Models
@@
    Long-Context MLLMs: We use state-of-the-art LLMs with long context windows to assess their zero-shot long-form writing capabilities. We use Qwen-Plus as the LLM to be compared in this paper, of which context window is 131,072 with maximum output of 8,192 tokens [^3] See the official website  https://www.alibabacloud.com/help/en/model-studio/what-is-qwen-llm for more details..
@@
Benchmark
We collect a bunch of annual reports spanning across Education, Trade, Health, Refugee and Climate domains, each ranges from several years. Each of them are public accessible, which offers a fair comparison between our proposed offline method and online search methods. The statistics of these five datasets are given in the Table 2. As the table shows, the real-world information are diverse and rich of multimodal information, which can offer valuable information.
Dataset # PDF # pages # avg pages # images
Education 1919 $8981$ $472\pm 72$ $6292$
Refugee 2020 $1979$ $98\pm 114$ $4644$
Climate $13$ $4764$ $366\pm 93$ $4232$
Health 2121 $2731$ $130\pm 33$ $2780$
Finance $22$ $5228$ $237\pm 71$ $2561$
Table 2: Statistics of evaluation datasets
To run Deepwriter, we manually download the five datasets, serving as the corpus $\mathcal{K}$, then we use MinerU to preprocess them and store the processed results in the database with the corresponding embeddings. Then, we ask offline methods to write a long article for a specific topic $Q$. For online methods, we use the online demo version to generate the corresponding article with the same topic $Q$.
Evaluation
To evaluate the quality of generated article, following (Shao et al. 2024; Jiang et al. 2024; Xi et al. 2025), we adopt Prometheus2-7B (Kim et al. 2024) as the judge model to evaluate the quality of the generate article from several dimensions including Interest Level; Coherence and Organization, Relevance and Focus; Coverage. The evaluation standard is given in Table 3.
Since Prometheus2-7B can only evaluate the text modality, we opt to use GPT-4o as the judge model to evaluate the coherence of each image and its surrounding paragraphs,
Evaluation Results
Performance
@@
Ablation study
We mainly discuss the impact of different modules to the final performance in this section.
Case Study
@@
Conclusion